Joe Thea and Taylor invested in a partnership in the ratio 1:4:7 respectively. Years later when the partnership is worth $1.6 million Thea decides to go to graduate school and sells her part of the partnership to Hoe. How much would Joe have to pay Thea to buy her share of the business?
Answer by Theo(13342)

the ratio is 1:4:7
the total number of units is 12.
joe has 1/12 of 1.6 million.
thea has 4/12 of 1.6 million.
taylor has 7/12 of 1.6 million.
thea's share comes out to be 1.6 million * 4/12 = .53333333.... million.
that comes out to be 533,333 and 1/3 dollars.
that's what joe would have to pay thea.

algebraically, you would solve as follows:

ratio is 1:4:7
let x = number of units that each one has.
total number of units is equal to 1x + 4x + 7x = 12x.
you get 12x = 6.1 million.
solve for x to get x = 6.1 million / 12.
joe has 1x = 1 * 6.1 million / 12.
thea has 4x = 4 * 6.1 million / 12.
taylor has 7x = 7 * 6.1 million /12.

you get the same answer either way.